OpenStudy (anonymous):

Find x if y=-2 4x-6/y =15 A.9/2 B.5 C.3 D.-6/23

OpenStudy (anonymous):

C. 3

OpenStudy (anonymous):

4*x-(6/(-2)) = 15 // - 15 4*x-(6/(-2))-15 = 0 4*x-15+3 = 0 4*x-12 = 0 // + 12 4*x = 12 // : 4 x = 12/4 x = 3